Dickson, Inc., has a debt-equity ratio of 2.4. The firm’s weighted average cost of capital is 9 percent and its pretax cost of debt is 7 percent. The tax rate is 25 percent. a. What is the company’s cost of equity capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. What is the company’s unlevered cost of equity capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) c. What would the company’s weighted average cost of capital be if the company's debt-equity ratio were .70 and 1.40? (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Answer a.

Debt-Equity Ratio = 2.40

Weight of Debt = 2.40 / 3.40
Weight of Equity = 1.00 / 3.40

WACC = Weight of Debt * Pretax Cost of Debt * (1 - Tax Rate) + Weight of Equity * Cost of Equity
0.09 = (2.40 / 3.40) * 0.07 * (1 - 0.25) + (1.00 / 3.40) * Cost of Equity
0.09 = 0.037059 + 0.294118 * Cost of Equity
0.294118 * Cost of Equity = 0.052941
Cost of Equity = 0.18 or 18.00%

Answer b.

Levered Cost of Equity = Unlevered Cost of Equity + (Unlevered Cost of Equity - Cost of Debt) * (1 - Tax Rate) * (Debt-Equity Ratio)
0.18 = Unlevered Cost of Equity + (Unlevered Cost of Equity - 0.07) * (1 - 0.25) * 2.40
0.18 = Unlevered Cost of Equity + 1.80 * Unlevered Cost of Equity - 0.1260
0.3060 = 2.80 * Unlevered Cost of Equity
Unlevered Cost of Equity = 0.1093 or 10.93%

Answer c.

If Debt-Equity Ratio is 0.70 is:

Levered Cost of Equity = Unlevered Cost of Equity + (Unlevered Cost of Equity - Cost of Debt) * (1 - Tax Rate) * (Debt-Equity Ratio)
Levered Cost of Equity = 0.1093 + (0.1093 - 0.07) * (1 - 0.25) * 0.70
Levered Cost of Equity = 0.1093 + 0.0206
Levered Cost of Equity = 0.1299 or 12.99%

WACC = Weight of Debt * Pretax Cost of Debt * (1 - Tax Rate) + Weight of Equity * Cost of Equity
WACC = (0.70 / 1.70) * 7.00% * (1 - 0.25) + (1.00 / 1.70) * 12.99%
WACC = 9.80%

If Debt-Equity Ratio is 1.40 is:

Levered Cost of Equity = Unlevered Cost of Equity + (Unlevered Cost of Equity - Cost of Debt) * (1 - Tax Rate) * (Debt-Equity Ratio)
Levered Cost of Equity = 0.1093 + (0.1093 - 0.07) * (1 - 0.25) * 1.40
Levered Cost of Equity = 0.1093 + 0.0413
Levered Cost of Equity = 0.1506 or 15.06%

WACC = Weight of Debt * Pretax Cost of Debt * (1 - Tax Rate) + Weight of Equity * Cost of Equity
WACC = (1.40 / 2.40) * 7.00% * (1 - 0.25) + (1.00 / 2.40) * 15.06%
WACC = 9.34%